OpenStudy (redneckoutlaw):

Marks: 4 [10.02] What are the solutions for x in the proportion 2 times x plus 3 all over x equals 2 times x all over 2.? Choose one answer. a. x = −1 and x = 3 b. x = 1 and x = −3 c. x = −2 and x = 3 d. x = 2 and x = −3 @wild_rebel_gurl21

OpenStudy (kirbykirby):

\[ \frac{2x+3}{x}=\frac{2x}{2} \, ?\]

OpenStudy (redneckoutlaw):

Yes, thats the question

OpenStudy (kirbykirby):

The right fraction can me simplified to just \(x\) since the 2's cancel out \[ \frac{2x+3}{x}=x\] cross multiply to get \[ 2x+3=x^2\\ x^2-2x-3=0\] can you factor this or use the quadratic formula to find the answer?

OpenStudy (anonymous):

(x+6) / 2x = x / 2 2(x+6) = 2x² -2x²+2x+12 = 0 divide both sides by -2 x²-x-6 = 0 (x-3)(x+2) = 0 x=3 x=-2

OpenStudy (anonymous):

so I thinik that the answer is C

OpenStudy (kirbykirby):

how did the numerator go from 2x+3 to x+6?

OpenStudy (redneckoutlaw):

0,0

OpenStudy (kirbykirby):

\( x^2-2x-3=0\\ (x-3)(x+1)=0\)

OpenStudy (redneckoutlaw):

So B

OpenStudy (kirbykirby):

not quite

OpenStudy (kirbykirby):

(x - 3) = 0, or (x + 1) = 0 x = 3, or x = -1

OpenStudy (anonymous):

no im telling you the answers C im correct

OpenStudy (anonymous):

but its ur testand ur call

OpenStudy (redneckoutlaw):

Wont load for me.

OpenStudy (kirbykirby):

well it also confirms that the solution is x=-1, x=3

OpenStudy (redneckoutlaw):

So A

OpenStudy (kirbykirby):

yea

OpenStudy (redneckoutlaw):

Ok bro thank you! :3mindhelping me with a few more?

OpenStudy (kirbykirby):

ok

OpenStudy (redneckoutlaw):

Thanks! Marks: 4 [10.03] Which expression is the simplified form of the square root of k to the seventeenth power.? Choose one answer. a. k8the square root of k. b. kthe square root of k to the fourth power. c. k4the square root of k. d. k5the square root of k to the seventh power.

OpenStudy (kirbykirby):

Recall that \(\Large \sqrt{x}=x^{\frac{1}{2}}\) So,\[\Large \sqrt{k^{17}}=\left( k^{17}\right)^{\frac{1}{2}}=k^{\frac{17}{2}}=k^{\frac{16}{2}+\frac{1}{2}}=k^8k^\frac{1}{2}=k^8\sqrt{k}\]

OpenStudy (redneckoutlaw):

So for the K 17/2 you minus 1 from the top?

OpenStudy (kirbykirby):

well I separated \(\large \frac{17}{2}=\frac{16+1}{2}=\frac{16}{2}+\frac{1}{2}\)

OpenStudy (redneckoutlaw):

Ahhhh ok

OpenStudy (redneckoutlaw):

Have time for another one?

OpenStudy (kirbykirby):

sure

OpenStudy (redneckoutlaw):

Marks: 4 [10.03] What is the simplified form of the square root of the quantity 72 times n to the thirteenth power.? Choose one answer. a. 8nthe square root of the quantity 3 times n to the sixth power. b. 3n6the square root of the quantity 8 times n. c. 8n6the square root of the quantity 3 times n. d. 6n6the square root of the quantity 2 times n.

OpenStudy (kirbykirby):

is that \[\Large \sqrt{72n^{13}} \]

OpenStudy (redneckoutlaw):

Yes

OpenStudy (redneckoutlaw):

Sorry if i take a couple to answer, had to help my dad

OpenStudy (kirbykirby):

\[\Large \sqrt{72n^{13}}\\~ \\\Large =\sqrt{72}\sqrt{n^{13}}\\ ~ \\ \Large =\sqrt{9\times 4 \times 2}\left(n^{13} \right)^{\frac{1}{2}}\\ ~ \\ \Large =\sqrt{9}\sqrt{4}\sqrt{2}\left(n^{\frac{13}{2}}\right) \\ ~ \\ \Large =3(2)\sqrt{2}\left( n^{\frac{12}{2}}n^{\frac{1}{2}}\right)\\ \Large =6n^6\sqrt{2n}\]

OpenStudy (redneckoutlaw):

Thanks man! Can u be like my tutor lol

OpenStudy (kirbykirby):

your welcome :) hehe well you can tag me in a question if you need it ;)

OpenStudy (redneckoutlaw):

Well i have a few more if your not to busy

OpenStudy (kirbykirby):

Can you post it in a new question? It's getting quite long the post here, and the more formulas posted, the more it lags :\